OpenStudy (anonymous):

The expression ( cot^2x/1-sin^2x is equivalent to ??

OpenStudy (tkhunny):

You have not written what you intend. Please remember your Order of Operations and use parentheses where necessary.

OpenStudy (anonymous):

cot^2x/ 1-sin^2x

OpenStudy (anonymous):

sorry there's no parentheses in the question

OpenStudy (tkhunny):

No, try cot^2x/ ( 1-sin^2x ). It is not the same in print as it is inline. You must see the difference and compensate.

OpenStudy (tkhunny):

Change cotangent to cos(x) / sin(x) and change 1-sin^2(x) to cos^2(x)

OpenStudy (anonymous):

what should I do after that ? please help

OpenStudy (tkhunny):

Did you do it? \(\dfrac{\cos^{2}(x)}{\sin^{2}(x)}\cdot \dfrac{1}{\cos^{2}(x)}\) Are you sure you can't come up with something?

OpenStudy (anonymous):

I got you :) Thank you sooo much

OpenStudy (anonymous):

I did it wrong the first time

OpenStudy (tkhunny):

Rule of thumb -- If you don't see it right away, change it to sine and cosine and hope...

OpenStudy (anonymous):

I'll remember that rule

OpenStudy (tkhunny):

Don't forget the "hope" part. There is always hope! :-)